# Revisiting Singlet Fermion Dark Matter with a Scalar Portal: Connecting Higgs Phenomenology and Strong Electroweak Phase Transition

### v2

@@ 1 Introduction
The Standard Model (SM) of particle physics, despite its remarkable success in describing the known fundamental particles and their interactions, remains incomplete. In particular, it fails to account for the observed baryon asymmetry of the Universe, namely the excess of matter over antimatter. Electroweak baryogenesis offers an attractive explanation for this asymmetry, provided that the electroweak phase transition (EWPT) is strongly first order. However, for the experimentally measured Higgs boson mass of $125\penalty 10000\ \mathrm{GeV}$, the electroweak phase transition in the SM proceeds as a smooth crossover Kajantie et al. (1996); Huet and Sather (1995); Csikor et al. (1999), precluding the strong out of equilibrium dynamics required to satisfy the Sakharov conditions Sakharov (1967). Consequently, electroweak baryogenesis is not viable within the SM framework Anderson and Hall (1992); Morrissey and Ramsey-Musolf (2012), motivating the exploration of physics beyond the SM. Among the many possible extensions, minimal and economical scenarios that can address these shortcomings without introducing unnecessary complexity are particularly appealing.
@@
Though SFOEWPT can be realized in a variety of new-physics scenarios, including models with extended scalar sectors such as two Higgs doublet models (2HDMs) Cline and Lemieux (1997); Bhatnagar et al. (2025), various singlet or multi-scalar extensions Díaz Sáez et al. (2021a); Ghorbani (2021); Ellis et al. (2023); Barger et al. (2008); Carena et al. (2020); Espinosa et al. (2012); Noble and Perelstein (2008); Cline and Kainulainen (2013); Alanne et al. (2014); Cline et al. (2013); Srivastava et al. (2025); Chaudhuri and Das (2022); Borah et al. (2025b); Chiang and Lu (2020); Kang et al. (2018); Kannike et al. (2020); Ghorbani (2024); Ghorbani and Ghorbani (2019); Díaz Sáez et al. (2024, 2021b); Murai et al. (2025); Chakrabarty et al. (2024); Ferber et al. (2024), or higher-dimensional operators Anderson and Hall (1992); Zhang (1993); Camargo-Molina et al. (2021); Hashino and Ueda (2023); Oikonomou and Giovanakis (2024); Gazi et al. (2024); we stick to a minimal extension of the SM. Thus, such minimal particle content beyond the SM within the renormalizable and gauge-invariant framework is not only able to explain the observed dark matter relic abundance, but also modifies the structure of the scalar potential in such a way that opens the possibility of a strong first-order phase transition, which is one of the required conditions for baryogenesis in the present universe. In addition, the generation of detectable stochastic gravitational waves from the expansion and subsequent collision of bubbles of the broken phase offers another experimental channel to test such a scenario Kamionkowski et al. (1994); Ellis et al. (2019); Croon et al. (2018); Beniwal et al. (2019); Huang and Zhang (2019); Hashino et al. (2018); Demidov et al. (2018); Mazumdar and White (2019); Kobakhidze et al. (2017b, a); Dev and Mazumdar (2016). The breakthrough discovery by LIGO Abbott and others (2016) has opened a new window into early universe dynamics, motivating the study of electroweak-scale GWs as a complementary probe of new physics in collider. The amplitude and peak frequency of the resulting GW spectrum encode information about the underlying scalar potential, making future space-based interferometers—such as LISA Seoane and others (2013), DECIGO Kawamura and others (2011), BBO Corbin and Cornish (2006), TAIJI Gong and others (2015), and TianQin Luo and others (2016)—promising tools for testing SFOEWPT scenarios. Therefore, the mixing angle between the SM doublet and real singlet that appears in the Higgs couplings controls the phenomenology of the heavier scalar state. On top of that, the new scalar interactions allow for non-resonant and resonant collider signatures - thus offering a well-defined and testable link between cosmological considerations and collider experiments.
@@ 3.3 LHC constraints on the heavy scalar $h_{2}$
At the High-Luminosity LHC, the di-Higgs sensitivity is expected to improve substantially 87. The projected limits strengthen the constraint on the mixing angle to roughly
@@ 4.1 Relic density
Figure 3: Variation of relic abundance with dark matter mass $m_{\chi}$ with two different chosen values of $\sin\theta$ and $g_{\chi}$ mentioned in the inset. Left panel: $m_{h_{2}}=200$ GeV, $\lambda_{hs}=2.0$, Right panel: $m_{h_{2}}=350$ GeV, $\lambda_{hs}=4.0$. The black dashed line indicates the observed dark matter relic abundance, $\Omega_{\rm DM}h^{2}=0.12\pm 0.001$ Aghanim and others (2020).
Figure 4: Variation of the Yukawa coupling $g_{\chi}$ with DM mass $m_{\chi}$ maintaining relic abundance in the range $0.1\leq\Omega_{\rm DM}h^{2}\leq 0.121$. The color gradient represents the mixing angle $\sin\theta$. Left for $m_{h_{2}}=200$ GeV and $\lambda_{hs}=$ 2.0. Right: $m_{h_{2}}=350$ GeV and $\lambda_{hs}=4.0$.
Overall, these patterns clearly indicate that achieving the observed thermal relic density requires the DM to be annihilated either through a sufficiently large mixing angle in the scalar sector or through a sizeable DM Yukawa interaction. We have displayed the relic abundance against the DM mass in Fig. 3 for some representative fixed values of the mixing angle $\sin\theta$ and $g_{\chi}$. The purpose is to highlight the dependency of the relic abundance on the mass of the DM. One may take, in principle, any mixing and coupling constant angle in between the representative points (i.e., $\sin\theta$ and $g_{\chi}$), but the observed relic density can be obtained only in the mentioned three regions of the DM mass.
The thermal relic abundance in this scalar-portal framework is governed by the interplay between the scalar mixing angle $\sin\theta$ and the dark-sector Yukawa coupling $g_{\chi}$. The left panel of Fig. 4 shows the allowed parameter space in the $(m_{\chi},\,g_{\chi})$ plane, where the color gradient encodes the value of the mixing angle $\sin\theta$, for a fixed heavy-scalar mass $m_{h_{2}}=200\penalty 10000\ \text{GeV}$, portal coupling $\lambda_{hs}=2.0$, and trilinear parameter $\mu_{3}=100\penalty 10000\ \text{GeV}$. The displayed region corresponds to parameter values that yield a relic density within the range, $0.1\leq\Omega_{\rm DM}h^{2}\leq 0.121$ Aghanim and others (2020).
@@
In contrast, when the dark-matter mass approaches the heavy-scalar resonance, $m_{\chi}\simeq m_{h_{2}}/2$, the Breit–Wigner enhancement substantially increases the annihilation rate. This resonant behaviour permits a wide range of mixing angles to satisfy the relic-density constraint, largely independent of the magnitude of $g_{\chi}$ (depending on how close the DM mass is to $m_{h_{2}}/2$). However, away from this near-degenerate region, the annihilation cross section becomes insufficient for small $g_{\chi}$, and even large values of $\sin\theta$ fail to reproduce the observed relic abundance. The right panel of Fig. 4, corresponding to a heavier scalar mass ($m_{h_{2}}=350\penalty 10000\ \text{GeV}$), is shown for the same trilinear coupling as in the left panel but with a larger Higgs-portal coupling, $\lambda_{hs}=4.0$. This choice of $\lambda_{hs}$ is motivated by the requirement of a strong first-order electroweak phase transition (FOEWPT) discussed in a later section. The qualitative behaviour remains similar to that observed in the left panel; however, the viable parameter space is shifted toward larger dark-matter masses, reflecting the dependence of the annihilation dynamics on the heavy-scalar mass scale.
@@
We have selected nine benchmark points (BPs) shown in Tab. 1 from various parts of the parameter space. The specified BPs remain viable under current experimental scrutiny, adhering to both cosmological relic density measurements and direct detection constraints. We perform our analysis for two representative heavy-scalar masses, $m_{h_{2}}=200$ and $350\penalty 10000\ \text{GeV}$, with portal couplings $\lambda_{hs}\sim 2.0$ and $4.0$, respectively. These mass choices ensure that the additional scalar remains within the LHC kinematic reach, allowing for potential production and providing an immediate experimental test of the model. For one of our benchmark scenarios (BP9), we consider $m_{h_{2}}$ to be less than the mass of the SM Higgs boson. This choice makes the selected set of benchmark points exhaustive, thereby rendering our systematic investigation complete, as it covers a wide region of parameter space that satisfies various collider bounds while simultaneously reproducing the correct DM and EWPT observables.
@@
Figure 5: Variation of mixing angle $\sin\theta$ with DM mass $m_{\chi}$ maintaining relic abundance in the range $0.1\leq\Omega_{\rm DM}h^{2}\leq 0.121$. The color gradient represents the Yukawa coupling $g_{\chi}$. Left for $m_{h_{2}}=200$ GeV and $\lambda_{hs}=$ 2.0. Right: $m_{h_{2}}=350$ GeV and $\lambda_{hs}=4.0$.
@@
Table 1: Benchmark points for singlet scalar extension of the SM along with singlet fermion dark matter. All the BPs satisfy the observed relic density and bound from upper limit of spin-independent DM-nucleon scattering cross section.
@@ 4.2 Direct detection
$\sigma_{\rm SI}=\frac{f_{N}^{2}m_{N}^{4}}{\pi(m_{\chi}+m_{N})^{2}}\left[g_{\chi}\,\frac{\sin\theta\cos\theta}{v}\left(\frac{1}{m_{h_{2}}^{2}}-\frac{1}{m_{h_{1}}^{2}}\right)\right]^{2}.$ (31)
@@
Figure 6: Variation of nucleon-dark matter spin independent scattering cross-section $\sigma_{\rm SI}$ with DM mass $m_{\chi}$ maintaining relic abundance in the range $0.1\leq\Omega_{\rm DM}h^{2}\leq 0.121$. The color gradient represents the mixing angle $\sin\theta$. Left for $m_{h_{2}}=200$ GeV and $\lambda_{hs}=$ 2.0. Right: $m_{h_{2}}=350$ GeV and $\lambda_{hs}=4.0$. The solid green line represents the latest $2\sigma$ bound from the LZ (2024) experiment Aalbers and others (2025).
@@
  • The bracket in Eq. (31) can become very small when the two propagators nearly cancel. This “destructive interference” or “blind-spot” behaviour produces the dip visible in Fig. 6 for certain combinations of $(m_{h_{2}},\sin\theta)$ and allows otherwise large-coupling scenarios to evade present bounds.
@@
  • In the region where the dark-matter mass satisfies $m_{\chi}\gtrapprox m_{h_{2}}$, the relic density is mostly insensitive to $\sin\theta$, but can be achieved only for relatively large values of $g_{\chi}$ (e.g., BP1 to BP6 and BP9). However, larger $g_{\chi}$ enhances spin independent direct detection cross section. Therefore, $\sigma_{\rm SI}$ permits only very small values of the mixing angle $\sin\theta$.
@@ 5.1 One-loop effective potential
The one-loop correction to the zero-temperature effective potential, known as the Coleman–Weinberg (CW) potential Coleman and Weinberg (1973), arises from the sum of all one-particle-irreducible vacuum diagrams and encodes quantum fluctuations around the classical background field. In dimensional regularization and within the $\overline{MS}$ scheme, its explicit form follows the standard expression Quiros (1999)
@@ 5.2 Electroweak phase transition
From the perspective of direct-detection constraints, the Higgs–singlet mixing angle must be very small, which in turn implies that the associated trilinear coupling $\mu_{hs}$ is highly suppressed. As a result, $\mu_{hs}$ has a subdominant role on the phase-transition dynamics. The impact of the singlet trilinear coupling $\mu_{3}$ on the EWPT in this setup can be summarized as follows. A sizable negative value of $\mu_{3}$ tends to enhance the barrier in the effective potential, thereby strengthening the phase transition. Consequently, for a fixed singlet mass, a relatively smaller Higgs portal coupling is sufficient to realize a SFOEWPT. In contrast, when $\mu_{3}$ is positive and large, its contribution weakens the transition, requiring a larger value of $\lambda_{hs}$ to get a SFOEWPT. We also find that a SFOEWPT can be achieved even when $\mu_{3}$ is set to zero. In the limit $\mu_{3}=0$ and $\mu_{hs}\to 0$, the model effectively reduces to a $\mathbb{Z}_{2}$-symmetric scalar extension, for which the correlation between $\lambda_{hs}$ and the singlet mass has been extensively discussed in Refs. Cline et al. (2013); Chaudhuri and Das (2022); Vaskonen (2017); Kang et al. (2018). It is worth noting, however, that Ref. Ellis et al. (2023) identifies an alternative scenario in which the explicit $\mathbb{Z}_{2}$-breaking parameters $\mu_{3}$ and $\mu_{hs}$ can play a significant role in strengthening the phase transition when the singlet acquires a nonzero VEV at zero temperature. Our results therefore indicate that the relevance of these parameters is strongly dependent on the underlying model and vacuum structure, and must be evaluated on a case-by-case basis.
Finally, we comment on the role of the BSM fermion $\chi$ in EWPT. Its contributions to the effective potential come from three sources. First, through thermal resummation, $\chi$ induces a temperature-dependent correction to the singlet scalar thermal mass via its Yukawa interaction (see Eq. (48)). In addition, $\chi$ contributes at one loop through both the CW correction and the finite-temperature part of the potential. These fermionic effects can strengthen the phase transition, with their impact controlled primarily by the mixed term $2g_{\chi}m_{\chi}s_{c}$ appearing in the field-dependent fermion mass squared evaluated at the critical temperature. When this combination becomes comparable to $T_{c}^{2}$, the fermionic contribution to the effective potential is maximized, thereby exerting its strongest influence on the phase-transition dynamics.

### v3

@@ 1 Introduction
The Standard Model (SM) of particle physics, despite its remarkable success in describing the known fundamental particles and their interactions, remains incomplete. In particular, it fails to account for the observed baryon asymmetry of the Universe, namely the excess of matter over antimatter. Electroweak baryogenesis offers an attractive explanation for this asymmetry, provided that the electroweak phase transition (EWPT) is strongly first order. However, for the experimentally measured Higgs boson mass of $125\penalty 10000\ \mathrm{GeV}$, the electroweak phase transition in the SM proceeds as a smooth crossover Kajantie et al. (1996); Huet and Sather (1995); Csikor et al. (1999), precluding the strong out of equilibrium dynamics required to satisfy the Sakharov conditions Sakharov (1967). Consequently, electroweak baryogenesis is not viable within the SM framework Anderson and Hall (1992); Morrissey and Ramsey-Musolf (2012), motivating the exploration of physics beyond the SM (BSM). Among the many possible extensions, minimal and economical scenarios that can address these shortcomings without introducing unnecessary complexity are particularly appealing.
@@
Though SFOEWPT can be realized in a variety of new-physics scenarios, including models with extended scalar sectors such as two Higgs doublet models (2HDMs) Cline and Lemieux (1997); Bhatnagar et al. (2025), various singlet or multi-scalar extensions Díaz Sáez et al. (2021a); Ghorbani (2021); Ellis et al. (2023); Ghorbani and Ghorbani (2020); Barger et al. (2008); Carena et al. (2020); Espinosa et al. (2012); Noble and Perelstein (2008); Cline and Kainulainen (2013); Alanne et al. (2014); Cline et al. (2013); Srivastava et al. (2025); Chaudhuri and Das (2022); Borah et al. (2025b); Chiang and Lu (2020); Kang et al. (2018); Kannike et al. (2020); Ghorbani (2024); Ghorbani and Ghorbani (2019); Díaz Sáez et al. (2024, 2021b); Murai et al. (2025); Chakrabarty et al. (2024); Ferber et al. (2024), or higher-dimensional operators Anderson and Hall (1992); Zhang (1993); Camargo-Molina et al. (2021); Hashino and Ueda (2023); Oikonomou and Giovanakis (2024); Gazi et al. (2024); we stick to a minimal extension of the SM. Thus, such minimal particle content beyond the SM within the renormalizable and gauge-invariant framework is not only able to explain the observed dark matter relic abundance, but also modifies the structure of the scalar potential in such a way that opens the possibility of a strong first-order phase transition, which is one of the required conditions for baryogenesis in the present universe. In addition, the generation of detectable stochastic gravitational waves from the expansion and subsequent collision of bubbles of the broken phase offers another experimental channel to test such a scenario Kamionkowski et al. (1994); Ellis et al. (2019); Croon et al. (2018); Beniwal et al. (2019); Huang and Zhang (2019); Hashino et al. (2018); Demidov et al. (2018); Mazumdar and White (2019); Kobakhidze et al. (2017b, a); Dev and Mazumdar (2016). The breakthrough discovery by LIGO Abbott and others (2016) has opened a new window into early universe dynamics, motivating the study of electroweak-scale GWs as a complementary probe of new physics in collider. The amplitude and peak frequency of the resulting GW spectrum encode information about the underlying scalar potential, making future space-based interferometers—such as LISA Seoane and others (2013), DECIGO Kawamura and others (2011), BBO Corbin and Cornish (2006), TAIJI Gong and others (2015), and TianQin Luo and others (2016)—promising tools for testing SFOEWPT scenarios. Therefore, the mixing angle between the SM doublet and real singlet that appears in the Higgs couplings controls the phenomenology of the heavier scalar state. On top of that, the new scalar interactions allow for non-resonant and resonant collider signatures - thus offering a well-defined and testable link between cosmological considerations and collider experiments.
@@ 3.3 LHC constraints on the heavy scalar $h_{2}$
At the High-Luminosity LHC, the di-Higgs sensitivity is expected to improve substantially 88. The projected limits strengthen the constraint on the mixing angle to roughly
@@ 4.1 Relic density
Figure 3: Variation of relic abundance with dark matter mass $m_{\chi}$ with two different chosen values of $\sin\theta$ and $g_{\chi}$ mentioned in the inset. Left panel for $m_{h_{2}}=200$ GeV and right for $m_{h_{2}}=350$ GeV. For both cases, we have taken $\lambda_{hs}=0.8$ and $\mu_{3}=-50$ GeV. The black dashed line indicates the observed dark matter relic abundance, $\Omega_{\rm DM}h^{2}=0.12\pm 0.001$ Aghanim and others (2020).
Figure 4: Variation of the Yukawa coupling $g_{\chi}$ with DM mass $m_{\chi}$ maintaining relic abundance in the range $0.1\leq\Omega_{\rm DM}h^{2}\leq 0.121$. The color gradient represents the mixing angle $\sin\theta$. Left panel for $m_{h_{2}}=200$ GeV and right for $m_{h_{2}}=350$ GeV. For both cases, we have taken $\lambda_{hs}=0.8$ and $\mu_{3}=-50$ GeV.
Overall, these patterns clearly indicate that achieving the observed thermal relic density requires the DM to be annihilated either through a sufficiently large mixing angle in the scalar sector or through a sizeable DM Yukawa interaction. We have displayed the relic abundance against the DM mass in Fig. 3 for some representative fixed values of the mixing angle $\sin\theta$ and coupling constant $g_{\chi}$. The purpose is to highlight the dependency of the relic abundance on the mass of the DM. One may take, in principle, any mixing angle and coupling constant in between the representative points, but the observed relic density can be obtained only in the three regions of the DM mass as mentioned above.
The thermal relic abundance in this scalar-portal framework is governed by the interplay between the scalar mixing angle $\sin\theta$ and the dark-sector Yukawa coupling $g_{\chi}$. The left panel of Fig. 4 shows the allowed parameter space in the $(m_{\chi},\,g_{\chi})$ plane, where the color gradient encodes the value of the mixing angle $\sin\theta$, for a fixed heavy-scalar mass $m_{h_{2}}=200\penalty 10000\ \text{GeV}$, portal coupling $\lambda_{hs}=0.8$, and trilinear parameter $\mu_{3}=-50\penalty 10000\ \text{GeV}$. The displayed region corresponds to parameter values that yield a relic density within the range, $0.1\leq\Omega_{\rm DM}h^{2}\leq 0.121$ Aghanim and others (2020).
@@
In contrast, when the dark-matter mass approaches the heavy-scalar resonance, $m_{\chi}\simeq m_{h_{2}}/2$, the Breit-Wigner enhancement substantially increases the annihilation rate. This resonant behavior permits a wide range of mixing angles to satisfy the relic density constraint, and largely independent of the magnitude of $g_{\chi}$ (depending on how close the DM mass is to $m_{h_{2}}/2$). However, away from this near-degenerate region, the annihilation cross section becomes insufficient for small $g_{\chi}$, and even large values of $\sin\theta$ fail to reproduce the observed relic abundance. The right panel of Fig. 4, corresponding to a heavier scalar mass ($m_{h_{2}}=350\penalty 10000\ \text{GeV}$), is shown for the same trilinear and Higgs portal couplings as in the left panel. The qualitative behavior remains similar to that observed in the left panel; however, the viable parameter space is shifted toward larger dark-matter masses, reflecting the dependence of the annihilation dynamics on the heavy scalar mass scale.
The requirement of a strong first-order electroweak phase transition, discussed in a later section, favors relatively large values of the portal coupling $\lambda_{hs}$ for a larger singlet scalar mass. Fig. 5 shows that the observed relic density can be accommodated even for such enhanced values of $\lambda_{hs}$. In particular, a larger portal coupling allows for a broader region of viable parameter space in the regime $m_{\chi}\lesssim m_{h_{2}}$. Moreover, as discussed earlier, variations in $\lambda_{hs}$ do not significantly impact the determination of the direct detection cross section.
Figure 5: Variation of the relic density for a sizable Higgs portal coupling $\lambda_{hs}=2.0$, while all other parameters are fixed to the same values as in the left panel of Fig. 4.
@@
We have selected nine benchmark points (BPs) shown in Tab. 1 from various parts of the parameter space. The specified BPs remain viable under current experimental scrutiny, adhering to both cosmological relic density measurements and direct detection constraints. We perform our analysis primarily for two representative BSM scalar masses, $m_{h_{2}}=200$ GeV and $350\penalty 10000\ \text{GeV}$. These mass choices ensure that the additional scalar remains within the LHC kinematic reach, allowing for potential production and providing an immediate experimental test of the model. For one of our benchmark scenarios (BP9), we consider $m_{h_{2}}$ to be less than the mass of the SM Higgs boson. This choice makes the selected set of benchmark points [^1] To achieve a SFOEWPT, several benchmarks require large quartic couplings ($\lambda_{hs}$ and $\lambda_{s}$). These values may trigger Landau poles at an intermediate scale, rendering the model as an effective field theory valid only up to a cutoff $\Lambda$. However, because this study targets electroweak-scale phenomena—such as relic density, direct detection, collider observables, and gravitational wave signals—this feature does not impact our conclusions. exhaustive, thereby rendering our systematic investigation complete, as it covers a wide region of parameter space that satisfies various collider bounds while simultaneously reproducing the correct DM and EWPT observables.
@@
Figure 6: Variation of mixing angle $\sin\theta$ with DM mass $m_{\chi}$ maintaining relic abundance in the range $0.1\leq\Omega_{\rm DM}h^{2}\leq 0.121$. The color gradient represents the Yukawa coupling $g_{\chi}$. Left panel for $m_{h_{2}}=200$ GeV and right for $m_{h_{2}}=350$ GeV. For both cases, we have taken $\lambda_{hs}=0.8$ and $\mu_{3}=-50$ GeV.
@@
Table 1: Benchmark points for singlet scalar extension of the SM along with singlet fermion dark matter. All the BPs satisfy the observed relic density and bound from the upper limit of spin-independent DM-nucleon scattering cross section.
@@ 4.2 Direct detection
$\sigma_{\rm SI}=\frac{f_{N}^{2}m_{N}^{4}}{\pi(m_{\chi}+m_{N})^{2}}\left[\frac{g_{\chi}\,\sin\theta\cos\theta}{v}\left(\frac{1}{m_{h_{2}}^{2}}-\frac{1}{m_{h_{1}}^{2}}\right)\right]^{2}.$ (31)
@@
Figure 7: Variation of nucleon-dark matter spin independent scattering cross-section $\sigma_{\rm SI}$ with DM mass $m_{\chi}$ maintaining relic abundance in the range $0.1\leq\Omega_{\rm DM}h^{2}\leq 0.121$. The color gradient represents the mixing angle $\sin\theta$. Left panel for $m_{h_{2}}=200$ GeV and right for $m_{h_{2}}=350$ GeV. For both cases, we have taken $\lambda_{hs}=0.8$ and $\mu_{3}=-50$ GeV. The solid green line represents the latest $2\sigma$ bound from the LZ (2024) experiment Aalbers and others (2025).
@@
  • The expression inside bracket in Eq. (31) can become very small when the two propagators nearly cancel. This “destructive interference” or “blind-spot” behaviour produces the dip visible in Fig. 6 for certain combinations of $(m_{h_{2}},\sin\theta)$ and allows otherwise large-coupling scenarios to evade present bounds.
@@
  • In the region where the dark-matter mass satisfies $m_{\chi}\gtrapprox m_{h_{2}}$, the relic density is mostly insensitive to $\sin\theta$, but can be achieved only for relatively large values of $g_{\chi}$ (e.g., BP1, BP2 BP5 and BP9). However, larger $g_{\chi}$ enhances spin independent direct detection cross section. Therefore, $\sigma_{\rm SI}$ permits only very small values of the mixing angle $\sin\theta$.
@@ 5.1 One-loop effective potential
The one-loop correction to the zero-temperature effective potential, known as the Coleman–Weinberg (CW) potential Coleman and Weinberg (1973), arises from the sum of all one-particle-irreducible vacuum diagrams and encodes quantum fluctuations around the classical background field. In $\overline{MS}$ scheme, its explicit form follows the standard expression Quiros (1999)
@@ 5.2 Electroweak phase transition
From the perspective of direct-detection constraints, the Higgs–singlet mixing angle must be very small, which in turn implies that the associated trilinear coupling $\mu_{hs}$ is highly suppressed. As a result, $\mu_{hs}$ has a subdominant role on the phase-transition dynamics. The impact of the singlet trilinear coupling $\mu_{3}$ on the EWPT in this setup can be summarized as follows. A sizable negative value of $\mu_{3}$ tends to enhance the barrier in the effective potential, thereby strengthening the phase transition. Consequently, for a fixed singlet mass, a relatively smaller Higgs portal coupling is sufficient to realize a SFOEWPT. In contrast, when $\mu_{3}$ is positive and large, its contribution weakens the transition, requiring a larger value of $\lambda_{hs}$ to get a SFOEWPT. We also find that a SFOEWPT can be achieved even when $\mu_{3}$ is set to zero. In the limit $\mu_{3}=0$ and $\mu_{hs}\to 0$, the model effectively reduces to a $\mathbb{Z}_{2}$-symmetric scalar extension, for which the correlation between $\lambda_{hs}$ and the singlet mass has been extensively discussed in Refs. Cline et al. (2013); Chaudhuri and Das (2022); Vaskonen (2017); Kang et al. (2018).
Finally, we comment on the role of the BSM fermion $\chi$ in EWPT. Its contributions to the effective potential come from three sources. First, through thermal resummation, $\chi$ induces a temperature-dependent correction to the singlet scalar thermal mass via its Yukawa interaction (see Eq. (48)). In addition, $\chi$ contributes at one loop through both the CW correction and the finite-temperature part of the potential. These fermionic contributions can affect the phase transition, with their impact controlled primarily by the mixed term $2g_{\chi}m_{\chi}s_{c}$ appearing in the field-dependent fermion mass squared evaluated at the critical temperature. When this combination becomes comparable to $T_{c}^{2}$, the fermionic contribution to the effective potential is maximized, thereby exerting its strongest influence on the phase-transition dynamics.
@@
Table 4: Phase transition parameters. All the VEVs and temperatures are in units of GeV.